OpenStudy (anonymous):

Find the length of the major axis for this ellipse: x^2+4y^2=33 Length of major axis= __square root of _____

OpenStudy (ash2326):

An ellipse is of the form \[\frac {x^2} {a^2} + \frac {y^2}{ b^2}=1\] 2a is the length of the major axis 2b is the length of the minor axis Can you do it now ?

OpenStudy (anonymous):

I know that, but they are asking it in a radical form, as is _____square root of ______ There is where I got lost...

OpenStudy (ash2326):

Can you convert the equation into the form of ellipse equation?

OpenStudy (anonymous):

yes, but what do I do with the 4y^2/33?

OpenStudy (ash2326):

We need 2a, don't bother about b^2!!

OpenStudy (anonymous):

The major axis is parallel to the x-axis.

OpenStudy (ash2326):

you could write in the form as \[\frac {x^2}{33}+\frac {y^2} {\frac {33}{4}}=1\] \[\large \frac {x^2}{\sqrt {33}}+\frac {y^2} {\sqrt{\frac {33}{4}}}=1\] \[\large \frac {x^2}{\sqrt {33}}+\frac {y^2} {{\frac {\sqrt{33}}{2}}}=1\]

OpenStudy (anonymous):

use the number under x^2. Since it is larger than the other one, it is a measure of the major axis. \[a=\sqrt{33}\]

OpenStudy (anonymous):

thank you guys :) I got it :D